# RecConv: Efficient Recursive Convolutions for Multi-Frequency Representations

### v1

@@ Abstract
Recent advances in vision transformers (ViTs) have demonstrated the advantage of global modeling capabilities, prompting widespread integration of large-kernel convolutions for enlarging the effective receptive field (ERF). However, the quadratic scaling of parameter count and computational complexity (FLOPs) with respect to kernel size poses significant efficiency and optimization challenges. This paper introduces RecConv, a recursive decomposition strategy that efficiently constructs multi-frequency representations using small-kernel convolutions. RecConv establishes a linear relationship between parameter growth and decomposing levels which determines the effective kernel size $k\times 2^{\ell}$ for a base kernel $k$ and $\ell$ levels of decomposition, while maintaining constant FLOPs regardless of the ERF expansion. Specifically, RecConv achieves a parameter expansion of only $\ell+2$ times and a maximum FLOPs increase of $5/3$ times, compared to the exponential growth ($4^{\ell}$) of standard and depthwise convolutions. RecNeXt-M3 outperforms RepViT-M1.1 by 1.9 $AP^{box}$ on COCO with similar FLOPs. This innovation provides a promising avenue towards designing efficient and compact networks across various modalities. Codes and models can be found at https://github.com/suous/RecNeXt.
@@ 1 Introduction
Figure 1: Effective Receptive Field (ERF) and Model Performance. (a) Comparison of effective kernel sizes and performance metrics among RecNeXt and prominent lightweight models. Top-1 accuracy is evaluated on ImageNet-1K and the latency is assessed using an iPhone 13 running iOS 18, where all models are formatted into mlmodel. $\dagger$” denotes pretraining at $256\times 256$ resolution and $\ast$” indicates pretraining with hard knowledge distillation. (b) The effective receptive fields of RepViT [64], FastViT [60], RepNeXt [72], and RecNeXt, where all models are pretrained with hard knowledge distillation. RecNeXt exhibits the broadest ERF, maintaining competitive latency and superior performance without reliance on structural reparameterization (SRP).
Convolutional Neural Networks (CNNs) [28, 27, 18, 49] have been foundational in computer vision for years, exploiting their intrinsic locality and translation equivariance [12]. Nonetheless, the recent emergence of vision transformers (ViTs) has demonstrated that establishing global context through self-attention mechanisms can lead to superior performance on various computer vision benchmarks [12, 2, 73]. Following the success of ViTs, extensive research has validated the effectiveness of large kernel convolutions in augmenting model capabilities through expanded receptive fields. ConvNeXt [38] and RepLKNet [9] achieved performance comparable or superior to Swin Transformers [37] when aligned with contemporary design paradigms. SLaK [36] and PeLK [3] explored kernels beyond $51\times 51$ and $101\times 101$ through dynamic sparsity and parameter sharing. Moreover, large-kernel CNNs have shown superiority as teachers in knowledge distillation for compact CNNs [24].
@@
OctConv [5] leveraged octave convolutions to enable multi-frequency feature representation. WTConv [13] incorporated small-kernel convolutions with cascade wavelet decomposition to obtain global receptive fields. While effective, their implementation requires substantial amount of parameters and resource-intensive operations. Motivated by the multi-frequency response observed in OctConv and WTConv, we propose RecConv — a recursively decomposed multi-frequency convolution optimized for resource-constrained mobile vision applications. Based on RecConv, we introduce RecNeXt, a simple yet effective vision backbone that delivers competitive efficiency with superior performance. Our contributions can be concluded as follows:
@@
  • Leveraging RecConv, we introduce RecNeXt, an efficient and effective vision backbone equipped with large receptive fields and optimized towards resource-constrained mobile vision applications.
  • Extensive experiments across various vision benchmarks demonstrate the superiority of RecNeXt over current leading approaches without relying on neural architecture search (NAS) or stractural reparameterization (SRP).
@@ 2 Related Work
The major perspective of our work is inspired by WTConv [13], while our design focuses on resource-constrained applications. And there are three major differences between WTConv and our proposed method: (1) We adopt a shared and learnable downsampling layer using a depthwise convolution with a stride of 2. (2) We maintain the same channel dimension across different scales for structural consistency and computational efficiency. (3) We use bilinear interpolation as an efficient upsampling operation to reduce parameter count and memory footprint.
@@ 3.1 Overall Architecture
Table 1: Classification performance on ImgeNet-1K. Following [32, 64], latency is measured on an iPhone 13 with models compiled by Core ML Tools. Similar to [14], throughput is tested on an Nvidia RTX3090 GPU with maximum power-of-two batch size that fits in memory. $\dagger$” denotes the evaluation image size is 256.
Model Type Params MACs Latency $\downarrow$ Throughput $\uparrow$ Top1 $\uparrow$ Auxiliary
(M) (G) (ms) (im/s) (%) NAS SRP
MobileViG-Ti [44] CNN-GNN 5.2 0.7 1.4 4337 75.7
SwiftFormer-XS [51] Hybrid 3.5 0.6 1.5 4304 75.7
EfficientFormerV2-S0 [34] Hybrid 3.5 0.4 1.4 1274 75.7
FastViT-T8$\dagger$ [60] Hybrid 3.6 0.7 1.3 3909 76.7
RepViT-M0.9 [64] CONV 5.1 0.8 1.3 4817 78.7
RepNeXt-M1 [72] CONV 4.8 0.8 1.3 3885 78.8
EfficientFormerV2-S1 [34] Hybrid 6.1 0.7 1.6 1153 79.0
RepViT-M1.0 [64] CONV 6.8 1.1 1.4 3910 80.0
RepNeXt-M2 [72] CONV 6.5 1.1 1.4 3198 80.1
RecNeXt-M1 CONV 5.2 0.9 1.4 384 79.2
RecNeXt-M2 CONV 6.8 1.2 1.5 325 80.3
MobileViG-S [44] CNN-GNN 7.2 1.0 1.6 2985 78.2
SwiftFormer-S [51] Hybrid 6.1 1.0 1.8 3376 78.5
EfficientFormer-L1 [32] Hybrid 12.3 1.3 1.8 3360 79.2
FastViT-T12$\dagger$ [60] Hybrid 6.8 1.4 1.8 3182 80.3
RepViT-M1.1 [64] CONV 8.2 1.3 1.5 3604 80.7
RepNeXt-M3 [72] CONV 7.8 1.3 1.5 2903 80.7
RecNeXt-M3 CONV 8.2 1.4 1.6 314 80.9
MobileViG-M [44] CNN-GNN 14.0 1.5 2.0 2491 80.6
FastViT-S12$\dagger$ [60] Hybrid 8.8 1.8 2.0 2313 80.9
SwiftFormer-L1 [51] Hybrid 12.1 1.6 2.2 2576 80.9
EfficientFormerV2-S2 [34] Hybrid 12.6 1.3 2.3 611 81.6
FastViT-SA12$\dagger$ [60] Hybrid 10.9 1.9 2.3 2181 81.9
RepViT-M1.5 [64] CONV 14.0 2.3 1.9 2151 82.3
RepNeXt-M4 [72] CONV 13.3 2.3 1.9 1745 82.3
RecNeXt-M4 CONV 14.1 2.4 2.4 169 82.5
EfficientFormer-L3 [32] Hybrid 31.3 3.9 3.4 1422 82.4
MobileViG-B [44] CNN-GNN 26.7 2.8 3.5 1446 82.6
SwiftFormer-L3 [51] Hybrid 28.5 4.0 3.4 1474 83.0
EfficientFormer-L7 [32] Hybrid 82.1 10.2 7.1 619 83.3
EfficientFormerV2-L [34] Hybrid 26.1 2.6 3.8 399 83.3
RepViT-M2.3 [64] CONV 22.9 4.5 2.8 1184 83.3
FastViT-SA24$\dagger$ [60] Hybrid 20.6 3.8 3.4 1128 83.4
RepNeXt-M5 [72] CONV 21.7 4.5 2.8 978 83.3
RecNeXt-M5 CONV 22.9 4.7 3.4 104 83.3
@@
The RecNeXt architecture builds upon RepViT [64] and RepNeXt [72], adhering to the four-stage framework of conventional CNNs [18] and hierarchical ViTs [37]. Starting with two $3\times 3$ convolutions with a stride of $2$ [64, 34, 67], progressively enhancing semantic representations while reducing spatial dimensions. The micro-blocks adopt the MetaNeXt design [38, 69], incorporating a token mixer for spatial feature extraction, a channel mixer for visual semantic interaction, a normalization layer [25] for training stability, and a shortcut connection [18] to facilitate smoother optimization [29].
@@ 3.2 Recursive Convolution
where $X_{\ell-1}\in\mathbb{R}^{B\times C\times 2h\times 2w}$, $X_{\ell}\in\mathbb{R}^{B\times C\times h\times w}$, and $X_{\ell+1}\in\mathbb{R}^{B\times C\times\frac{h}{2}\times\frac{w}{2}}$ represent feature maps at levels $\ell-1$, $\ell$, and $\ell+1$, respectively. $\mathrm{DWConv_{k\times k,\downarrow}}$ indicates the shared-weight depthwise convolution with a kernel size of $k\times k$ and a stride of $2$, while $\mathrm{DWConv_{k\times k}^{(\ell+1)}}$ denotes the depthwise convolution with a kernel size of $k\times k$ at level $\ell+1$. $\mathrm{Interp_{\uparrow}}$ stands for the bilinear interpolation upsampling operation. Features at different scales are progressively fused back into the original feature map, effectively replicating multi-frequency representations with large-kernel receptive fields.
Table 2: Complexity of different types of convolution. The measurement is streamlined by focusing on convolution operations, with the assumption of consistent input and output channels, and excluding the bias term. $k$, $C$, $H$ and $W$ denote the base kernel size, channel number, image height and width, respectively. And $\ell$ represents the number of decomposition levels that determine the effective kernel size, which is $k\times 2^{\ell}$.
@@
In addition, our RecConv embodies a versatile design philosophy that accommodates different operators and dimensions. For instance, the shared downsampling layer can be substituted with a sequence of channel-wise chunking operations followed by depthwise convolutions; depthwise convolutions at each level can be replaced by efficient attention mechanisms; bilinear upsampling operations can be converted to transposed convolutions; and element-wise addition can be reconfigured as channel-wise concatenation.
@@ 4.1 Image Classification
Table 3: Results without distillation on ImageNet-1K, where“$\dagger$” denotes the evaluation image size is 256.
@@
As demonstrated in Table 1, RecNeXt consistently delivers superior performance and competitive latency across various model sizes without requiring additional training auxiliaries. Notably, smaller RecNeXt models outperform their RepNeXt counterparts by 0.2% in top-1 accuracy while maintaining comparable latency and FLOPs. RecNeXt-M5 matches the top-1 accuracy (83.3%) of EfficientFormerV2-L with fewer parameters and lower latency. Without leveraging structural reparameterization (SRP) or neural architecture search (NAS), the RecNeXt series consistently surpasses SwiftFormer in top-1 accuracy with similar computational complexity. However, RecNeXt encounters challenges with throughput due to the extensive employment of bilinear upsample operations, which are less friendly to GPU acceleration. This deficiency can be alleviated by leveraging transposed depthwise convolutions at the cost of increased computational complexity and parameter count.
@@
Table 3 illustrates that RecNeXt maintains strong performance across various model scales. RecNeXt-M1 achieves a top-1 accuracy of 78.0% with a latency of 1.4 ms and 5.2M parameters, outperforming prior lightweight models such as RepViT-M0.9 (77.4%) and RepNeXt-M1 (77.5%) at comparable latencies. RecNeXt-M2 further improves to 79.2% accuracy with 6.8M parameters and 1.5 ms latency, surpassing RepViT-M1.0 (78.6%) and RepNeXt-M2 (78.9%) under similar resource constraints. For higher-capacity models, RecNeXt-M3 achieves a top-1 accuracy of 79.6%, exceeding other leading models without relying on structural reparameterization (SRP) and neural architecture search (NAS). However, larger models like RecNeXt-M4 and RecNeXt-M5 exhibit performance saturation alongside increased latency, with RecNeXt-M5 plateaus at 81.6% top-1 accuracy but with a latency of 3.4ms (0.6ms higher than RepNext-M5). Although RecNeXt-M5 experiences performance degradation, it still outperforms PoolFormer-S36 (81.4%) under the same training protocol, with fewer parameters and lower latency. Additionally, incorporating SRP or advanced operators could potentially mitigate this limitation.

### v2

@@ Abstract
Recent advances in vision transformers (ViTs) have demonstrated the advantage of global modeling capabilities, prompting widespread integration of large-kernel convolutions for enlarging the effective receptive field (ERF). However, the quadratic scaling of parameter count and computational complexity (FLOPs) with respect to kernel size poses significant efficiency and optimization challenges. This paper introduces RecConv, a recursive decomposition strategy that efficiently constructs multi-frequency representations using small-kernel convolutions. RecConv establishes a linear relationship between parameter growth and decomposing levels which determines the effective receptive field $k\times 2^{\ell}$ for a base kernel $k$ and $\ell$ levels of decomposition, while maintaining constant FLOPs regardless of the ERF expansion. Specifically, RecConv achieves a parameter expansion of only $\ell+2$ times and a maximum FLOPs increase of $5/3$ times, compared to the exponential growth ($4^{\ell}$) of standard and depthwise convolutions. RecNeXt-M3 outperforms RepViT-M1.1 by 1.9 $AP^{box}$ on COCO with similar FLOPs. This innovation provides a promising avenue towards designing efficient and compact networks across various modalities. Codes and models can be found at https://github.com/suous/RecNeXt.
@@ 1 Introduction
Figure 1: Effective Receptive Field (ERF) and Model Performance. (a) Comparison of ERFs and performance metrics among RecNeXt and prominent lightweight models. Top-1 accuracy is evaluated on ImageNet-1K and the latency is assessed using an iPhone 13 running iOS 18, where all models are formatted into mlmodel. $\dagger$” denotes pretraining at $256\times 256$ resolution and $\ast$” indicates pretraining with hard knowledge distillation. (b) The effective receptive fields of RepViT [69], FastViT [65], RepNeXt [78], and RecNeXt. RecNeXt exhibits the broadest ERF, maintaining competitive latency and superior performance without reliance on structural reparameterization (SRP).
Convolutional Neural Networks (CNNs) [32, 31, 21, 54] have been foundational in computer vision for years, exploiting their intrinsic locality and translation equivariance [13]. Nonetheless, the recent emergence of vision transformers (ViTs) has demonstrated that establishing global context through self-attention mechanisms can lead to superior performance on various computer vision benchmarks [13, 3, 79]. Following the success of ViTs, extensive research has validated the effectiveness of large kernel convolutions in augmenting model capabilities through expanded receptive fields. ConvNeXt [42] and RepLKNet [10] achieved performance comparable or superior to Swin Transformers [41] when aligned with contemporary design paradigms. SLaK [40] and PeLK [4] explored kernels beyond $51\times 51$ and $101\times 101$ through dynamic sparsity and parameter sharing. Moreover, large-kernel CNNs have shown superiority as teachers in knowledge distillation [27].
@@
OctConv [6] leveraged octave convolutions to enable multi-frequency feature representation. WaveMix [29] and WTConv [14] incorporated small-kernel convolutions with wavelet decomposition to obtain global receptive fields. HorNet [53] performed high-order spatial interactions with gated convolutions and recursive designs. While effective, their implementation requires substantial amount of parameters and resource-intensive operations. Motivated by their multi-frequency response, we propose a recursively decomposed convolution (RecConv) optimized for resource-constrained mobile vision applications. Based on RecConv, we introduce RecNeXt, a simple yet effective vision backbone that delivers competitive efficiency with superior performance. Our contributions can be concluded as follows:
@@
  • Leveraging RecConv, we introduce RecNeXt, two series of efficient and effective vision backbones equipped with large receptive fields and optimized towards resource-constrained mobile vision applications.
  • Extensive experiments across various vision benchmarks demonstrate the flexibility and effectiveness of RecNeXt without relying on neural architecture search (NAS) or structural reparameterization (SRP).
@@ 2 Related Work
The major perspective of our work is inspired by WTConv [14], while our design focuses on resource-constrained applications. And there are four major differences between WTConv and our proposed method: (1) We adopt a shared and learnable depthwise convolutional downsampling layer with a stride of 2. (2) We maintain the same channel dimension across different scales for structural consistency and computational efficiency. (3) We use bilinear interpolation (interchangeable during inference) to reduce parameter count and memory footprint. (4) Our design supports simultaneous recursive decomposition along both spatial and channel dimensions.
@@ 3.1 Overall Architecture
Table 1: Classification performance on ImageNet-1K. Following [36, 69], NPU latency is measured on an iPhone 13 with models compiled by Core ML Tools. CPU latency is accessed on a Quad-core ARM Cortex-A57 processor. Throughput is tested on an Nvidia RTX3090 GPU with maximum power-of-two batch size that fits in memory. $\dagger$” denotes the evaluation image size is 256. Gray annotations indicate replacing bilinear interpolation with nearest interpolation during inference, improving GPU throughput at a slight accuracy cost.
Model Type Params MACs Latency (ms) $\downarrow$ Throughput $\uparrow$ Top1 $\uparrow$ Auxiliary
(M) (G) NPU CPU (im/s) (%) NAS SRP
MobileViG-Ti [48] CNN-GNN 5.2 0.7 1.4 261 4337 75.7
SwiftFormer-XS [56] Hybrid 3.5 0.6 1.5 194 4304 75.7
EfficientFormerV2-S0 [38] Hybrid 3.5 0.4 1.4 198 1274 75.7
FastViT-T8$\dagger$ [65] Hybrid 3.6 0.7 1.3 280 3909 76.7
RepViT-M0.9 [69] CONV 5.1 0.8 1.3 222 4817 78.7
RepNeXt-M1 [78] CONV 4.8 0.8 1.3 267 3885 78.8
EfficientFormerV2-S1 [38] Hybrid 6.1 0.7 1.6 303 1153 79.0
RepViT-M1.0 [69] CONV 6.8 1.1 1.4 280 3910 80.0
RepNeXt-M2 [78] CONV 6.5 1.1 1.4 325 3198 80.1
RecNeXt-M1 CONV 5.2 0.9 1.4 361 384 / 3294 79.2 / 78.8
RecNeXt-M2 CONV 6.8 1.2 1.5 431 325 / 2707 80.3 / 80.1
MobileViG-S [48] CNN-GNN 7.2 1.0 1.6 383 2985 78.2
SwiftFormer-S [56] Hybrid 6.1 1.0 1.8 278 3376 78.5
EfficientFormer-L1 [36] Hybrid 12.3 1.3 1.8 299 3360 79.2
FastViT-T12$\dagger$ [65] Hybrid 6.8 1.4 1.8 456 3182 80.3
RepViT-M1.1 [69] CONV 8.2 1.3 1.5 326 3604 80.7
RepNeXt-M3 [78] CONV 7.8 1.3 1.5 373 2903 80.7
RecNeXt-M3 CONV 8.2 1.4 1.6 482 314 / 2491 80.9 / 80.5
MobileViG-M [48] CNN-GNN 14.0 1.5 2.0 488 2491 80.6
FastViT-S12$\dagger$ [65] Hybrid 8.8 1.8 2.0 549 2313 80.9
SwiftFormer-L1 [56] Hybrid 12.1 1.6 2.2 411 2576 80.9
EfficientFormerV2-S2 [38] Hybrid 12.6 1.3 2.3 498 611 81.6
FastViT-SA12$\dagger$ [65] Hybrid 10.9 1.9 2.3 569 2181 81.9
RepViT-M1.5 [69] CONV 14.0 2.3 1.9 540 2151 82.3
RepNeXt-M4 [78] CONV 13.3 2.3 1.9 632 1745 82.3
RecNeXt-M4 CONV 14.1 2.4 2.4 843 169 / 1476 82.5 / 82.0
EfficientFormer-L3 [36] Hybrid 31.3 3.9 3.4 795 1422 82.4
MobileViG-B [48] CNN-GNN 26.7 2.8 3.5 865 1446 82.6
SwiftFormer-L3 [56] Hybrid 28.5 4.0 3.4 859 1474 83.0
EfficientFormer-L7 [36] Hybrid 82.1 10.2 7.1 1906 619 83.3
EfficientFormerV2-L [38] Hybrid 26.1 2.6 3.8 877 399 83.3
RepViT-M2.3 [69] CONV 22.9 4.5 2.8 1007 1184 83.3
FastViT-SA24$\dagger$ [65] Hybrid 20.6 3.8 3.4 1044 1128 83.4
RepNeXt-M5 [78] CONV 21.7 4.5 2.8 1144 978 83.3
RecNeXt-M5 CONV 22.9 4.7 3.4 1487 104 / 828 83.3 / 83.0
@@
The RecNeXt architecture builds upon RepViT [69] and RepNeXt [78], adhering to the four-stage framework of conventional CNNs [21] and hierarchical ViTs [41]. Starting with two $3\times 3$ convolutions with a stride of $2$ [69, 38, 73], progressively enhancing semantic representations while reducing spatial dimensions. The micro-blocks adopt the MetaNeXt design [42, 75], incorporating a token mixer for spatial feature extraction, a channel mixer for visual semantic interaction, a normalization layer [28, 2] for training stability, and a shortcut connection [21] to facilitate smoother optimization [33].
@@ 3.2 Recursive Convolution
where $X_{\ell-1}\in\mathbb{R}^{B\times C\times 2h\times 2w}$, $X_{\ell}\in\mathbb{R}^{B\times C\times h\times w}$, and $X_{\ell+1}\in\mathbb{R}^{B\times C\times\frac{h}{2}\times\frac{w}{2}}$ represent feature maps at levels $\ell-1$, $\ell$, and $\ell+1$, respectively. $\mathrm{DWConv_{k\times k,\downarrow}}$ indicates the shared-weight depthwise convolution with a kernel size of $k\times k$ and a stride of $2$, while $\mathrm{DWConv_{k\times k}^{(\ell+1)}}$ denotes the depthwise convolution with a kernel size of $k\times k$ at level $\ell+1$. $\mathrm{Interp_{\uparrow}}$ stands for bilinear interpolation (interchangeable during inference) upsampling operation. Features at different scales are progressively fused back into the original feature map, effectively replicating multi-frequency representations with large-kernel receptive fields.
Table 2: Complexity of different types of convolution. The measurement is streamlined by focusing on convolution operations, with the assumption of consistent input and output channels, and excluding the bias term. $k$, $C$, $H$ and $W$ denote the base kernel size, channel number, image height and width, respectively. And $\ell$ represents the number of decomposition levels that determine the effective receptive field, which is $k\times 2^{\ell}$.
@@
In addition, our RecConv embodies a versatile design philosophy that accommodates different operators and dimensions. For instance, the shared downsampling layer can be substituted with a sequence of channel-wise chunking followed by depthwise convolutions; depthwise convolutions at each level can be replaced by efficient attention mechanisms [17, 18]; bilinear upsampling operations can be converted to transposed convolutions; element-wise addition can be reconfigured as channel-wise concatenation.
@@ 4.1 Image Classification
Table 3: Results on ImageNet-1K without distillation, where“$\dagger$” denotes evaluation at $256\times 256$ image size. RecNeXt achieves higher accuracy on smaller models, but experiences performance saturation with larger variants.
@@
As demonstrated in Table 1, RecNeXt consistently delivers superior performance and competitive latency across various model sizes without requiring additional training auxiliaries. Notably, smaller RecNeXt models outperform their RepNeXt counterparts by 0.2% in top-1 accuracy while maintaining comparable latency and FLOPs. RecNeXt-M5 matches the top-1 accuracy (83.3%) of EfficientFormerV2-L with fewer parameters and lower latency. Without leveraging structural reparameterization (SRP) or neural architecture search (NAS), the RecNeXt series consistently surpasses SwiftFormer in top-1 accuracy with similar computational complexity. However, RecNeXt encounters challenges with throughput due to the extensive employment of bilinear upsampling operations, which can be replaced by nearest interpolation during inference. And can also be alleviated by leveraging transposed depthwise convolutions at the cost of increased computational complexity and parameter count.
@@
Table 3 illustrates that RecNeXt maintains strong performance across various model scales. RecNeXt-M1 achieves a top-1 accuracy of 78.0% with a latency of 1.4 ms and 5.2M parameters, outperforming prior lightweight models such as RepViT-M0.9 (77.4%) and RepNeXt-M1 (77.5%) at comparable latencies. RecNeXt-M2 further improves to 79.2% accuracy with 6.8M parameters and 1.5 ms latency, surpassing RepViT-M1.0 (78.6%) and RepNeXt-M2 (78.9%) with similar FLOPs. For higher-capacity models, RecNeXt-M3 achieves a top-1 accuracy of 79.6%, exceeding other leading models without relying on structural reparameterization (SRP) and neural architecture search (NAS). However, larger models like RecNeXt-M4 and RecNeXt-M5 exhibit performance saturation alongside increased latency, with RecNeXt-M5 plateaus at 81.6% top-1 accuracy but with a latency of 3.4ms (0.6ms higher than RepNext-M5). Although RecNeXt-M5 experiences performance degradation, it still outperforms PoolFormer-S36 (81.4%) with fewer parameters and lower latency. Additionally, incorporating SRP, carefully tuned configurations, or advanced operators could potentially mitigate this limitation.